# Bridging Wright-Fisher and Moran models

### v1

@@ Abstract
The Wright-Fisher model and the Moran model are both widely used in population genetics. They describe the time evolution of the frequency of an allele in a well-mixed population with fixed size. We propose a simple and tractable model which bridges the Wright-Fisher and the Moran descriptions. We assume that a fixed fraction of the population is updated at each discrete time step. In this model, we determine the fixation probability of a mutant in the diffusion approximation, as well as the effective population size. We generalize our model, first by taking into account fluctuating updated fractions or individual lifetimes, and then by incorporating selection on the lifetime as well as on the reproductive fitness.
@@ I Introduction
Several population genetics models that are more general than the Wright-Fisher and Moran model have been developed. They include the Karlin and McGregor model [15, 16] and the Chia and Watterson model [17]. The latter considers a haploid population of fixed size $N$, where each individual reproduces independently, according to a specific distribution of offspring number. Then $r$ parents and $N-r$ offspring constitute the next generation, where $r$ is a random variable conditioned by the offspring population size. While this encompasses Wright-Fisher and Moran models as particular cases, Chia and Watterson [17] indicated that “there did not seem to be any continuity […]” between these two frameworks. Besides, as pointed out by Cannings [18], the Chia and Watterson model is rather complex since it involves the offspring distribution, the fraction of surviving offsprings, and the sampling of the new generation from the parent and the offspring populations. It makes it hardly tractable, and to our knowledge, the mutant fixation probability has not been derived under this model.
Here, we propose a simple and tractable model which bridges the Wright-Fisher model and the Moran model. Specifically, we assume that a fixed fraction of the population is updated at each discrete time step. We show that our model generalizes over both the Wright-Fisher model and the Moran model. Under Kimura’s diffusion approximation, we obtain a simple expression of the fixation probability of a mutant in our model. We further study the rate of approach to the steady state in the neutral case, and determine the variance effective population size and the eigenvalue effective population size. Next, we generalize our model, first by taking into account fluctuations of the fraction that is updated at each discrete time step, and then by considering individuals with fluctuating lifetimes. This allows us to express the fixation probability as a function of the mean life time of individuals. Finally, we further generalize our model by incorporating selection on the lifetime as well as on the reproductive fitness, i.e., both on birth and death processes, assuming that lifetimes are geometrically distributed.
@@ II.1 Description of the model
Link to the Wright-Fisher model.
@@
Link to the Moran model.
@@ II.2 Mutant fixation probability
Let us consider a large-sized population $N\gg 1$ and assume that $s$ is of order $1/N$. These conditions allow us to employ Kimura’s diffusion approximation [19] (see Supplemental Material (SM) for mathematical details [20]). The fixation probability $\phi(x)$ of the mutant type, starting from a fraction $x$ of mutants in the population, then satisfies the following Kolmogorov backward equation [19]:
@@
where $\Delta x$ is the variation of the mutant fraction occurring in the first discrete time step starting from the initial fraction $x$, and $\mathbb{E}(\Delta x)$ and $\mathbb{V}(\Delta x)$ denote respectively its expected value and variance. To leading order in $1/N$, they read (see SM [20] for details)
@@
This result generalizes over the fixation probability of both the Moran and the Wright-Fisher models in the diffusion regime. We recover them by taking $\rho=1/N$ (which is then neglected since we restrict to leading order in $1/N$), and $\rho=1$, respectively. In Fig. 2, we show the probability $\phi_{1}=\phi(1/N)$ of fixation in Eq. 11 for different values of the updated fraction $\rho$.
@@
We note that this is analogous to Haldane’s formula [21]
@@
where $\sigma^{2}=2-\rho$ is the variance of the number of offspring drawn from a Poissonian distribution (see SM [20] for details).
II.3 Rate of approach to steady state in the neutral case
In this section, we consider neutral mutants, i.e. $s=0$. The Markov chain representing the time evolution of the number of mutants in the population possesses two absorbing states, namely fixation and extinction of the mutant type. Therefore, the transition matrix in our model, whose elements are given by Eq. 4, has two unit eigenvalues [22]. The next largest non-unit eigenvalue provides information on the rate of decrease of genetic diversity in the population. It is thus of interest to calculate it. Denoting by $X_{t}$ be the number of mutant at time $t$, we can show that (see SM [20] for details)
@@ II.3 Rate of approach to steady state in the neutral case
The convergence to steady state can be quantified by considering the vector $\bm{\pi}_{t}=(p_{0}(t),p_{1}(t),p_{2}(t),...,p_{N}(t))$ where $p_{k}(t)$ denotes the probability of having $k$ mutants at time $t$ starting from $1$ mutant at $t=0$. For $t\rightarrow\infty$, we have (see SM [20])
@@
where ||.$||.||_{2}$ refers to the Euclidean norm and $\bm{\pi}_{\infty}=\lim\limits_{t\rightarrow\infty}\bm{\pi}_{t}=(1-1/N,0,0,...,0,1/N)$. In Fig. 3, this prediction is tested numerically for different updated fractions $\rho$ of the population. We find a good agreement between our simulation results and the long-time asymptotic behavior given in Eq. 17.
@@
Figure 3: Approach to steady state in the neutral case. We show the Euclidean norm of $\bm{\pi}_{t}-\bm{\pi}_{\infty}$ (see main text) versus time in generations (i.e. rescaled with the generation time $1/\rho$), for $N=1000$ and with different updated fractions $\rho$. Full lines: simulation results obtained from $5.10^{7}$ runs. Dashed lines: asymptotic behavior given in Eq. 17, with a multiplicative constant chosen for legibility.
II.4 Effective population sizes
@@
The effective population size is the size a regular Wright-Fisher population should have to match a given property of the population considered. It is of interest to express it for a population of size $N$ within our model.
Let us first consider the variance effective population size. Because the average generation time is $1/\rho$, the variance accumulated over one generation is $1/\rho$ times that in Eq. 10. Thus, the variance effective population size [3] within our model reads
$N_{e}=\frac{x(1-x)}{\mathbb{V}(\Delta x)/\rho}=\frac{N}{2-\rho}.$ (18)
@@
Next, let us consider the eigenvalue effective population size. Eq. 16 allows us to express it as [3]
$N_{e}=\frac{1}{1-r}=\frac{N}{2-\rho}\,.$ (19)
We observe that these two effective population sizes coincide.
@@ III.1 Fluctuating updated fraction
where $P_{i\rightarrow j}^{(N,M)}$ is given by Eq. 4. To obtain the mutant fixation probability in the diffusion approximation within this new model, we express the first two moments of $\Delta x$. The calculation is similar to the one presented in SM [20] for the model with fixed updated fraction, and yields to leading order in $1/N$
@@
[…]
@@
[…]
@@
$T\right>$ $=\frac{1}{\left<\rho\right>}.$ (27)
@@ III.2 Fluctuating lifetimes
$\lim\limits_{t\rightarrow+\infty}p_{u}(t)=\frac{1}{\langle T\rangle}\,.$ (29)
@@
$\langle\rho\rangle=\frac{1}{\langle T\rangle}\,,$ (30)
@@
$\frac{\langle\rho^{2}\rangle}{\langle\rho\rangle}=\frac{1}{\langle T\rangle}\,.$ (31)
@@ III.3 Selection on the lifetime
Assuming as usual that $s$ is of order $1/N$, we obtain, to leading order in $1/N$ (see SM [20]):
@@ IV Discussion
We proposed a simple and tractable model that bridges the Wright-Fisher model and the Moran model. In our model, a fraction $\rho$ of the population is updated at each discrete time step, using binomial sampling. The Wright-Fisher model and the Moran model are recovered where the updated fraction $\rho$ is respectively $1$ and $1/N$. Under the diffusion approximation, we obtained the mutant fixation probability, as well as the rate of approach to steady state and the variance and eigenvalue effective population sizes. We further extended our model to the case where the updated fraction $\rho$ fluctuates. We derived a simple relation between the mean updated fraction and the mean individual lifetime. We considered the case where the lifetime fluctuates. Finally, we examined the situation where selection occurs both on birth and on death, assuming that the lifetime of individuals is geometrically distributed. In all these cases, we obtained simple expressions of the fixation probability in the diffusion approximation.
Our model can be seen as a particular case of the Chia and Watterson model [17]. We do not focus on the offspring distribution, but on the fraction of the population that is updated at each time step. This leads to a more tractable model where the Wright-Fisher and the Moran model emerge as simple limiting cases in the fixation probabilities.
Considering intermediate updated fractions has concrete applications, including bet-hedging strategies, where only a fraction of the population gives rise to the next generation, while the other seeds or offspring-generating individuals are dormant. These strategies allow the population to survive unpredictable or harsh environment changes. They are common in annual plants [26, 27], bacterial populations [28], and insects [29].
@@
Here, we assumed discrete time steps, as in most classic models of population genetics. This is relevant in many situations, e.g. annual reproduction. Moreover, when continuous time is more relevant, as in bacterial reproduction, mutant replications remain synchronized for some time, as they stem from the same ancestor [30]. Besides, we focused on a well-mixed population with fixed size. It would be interesting to extend our model to spatially structured populations, or to fluctuations of the population size [31].
Supplemental Material for “Bridging Moran and Wright-Fisher models”

### v2

@@ Abstract
The Wright-Fisher model and the Moran model are both widely used in population genetics. They describe the time evolution of the frequency of an allele in a well-mixed population with fixed size. We propose a simple and tractable model which bridges the Wright-Fisher and the Moran descriptions. We assume that a fixed fraction of the population is updated at each discrete time step. In this model, we determine the fixation probability of a mutant and its average fixation and extinction times, under the diffusion approximation. We further study the associated coalescent process, which converges to Kingman’s coalescent, and we calculate effective population sizes. We generalize our model, first by taking into account fluctuating updated fractions or individual lifetimes, and then by incorporating selection on the lifetime as well as on the reproductive fitness.
@@ I Introduction
Several population genetics models that are more general than the Wright-Fisher and Moran model have been developed. They include the Karlin and McGregor model [15, 16], the Chia and Watterson model [17], and in the neutral case, the Cannings model [18]. The Chia and Watterson model considers a haploid population of fixed size $N$, where each individual reproduces independently, according to a specific distribution of offspring number. Then $r$ parents and $N-r$ offspring constitute the next generation, where $r$ is a random variable conditioned by the offspring population size. This encompasses Wright-Fisher and Moran models as particular cases. However, as pointed out by Cannings [19], the Chia and Watterson model is rather complex since it involves the offspring distribution, the fraction of surviving offsprings, and the sampling of the new generation from the parent and the offspring populations. It makes it hardly tractable, and to our knowledge, the mutant fixation probability has not been derived under this model.
Here, we consider a simple and tractable case of the Chia and Watterson model, which bridges Wright-Fisher and Moran models. Specifically, we assume that a fraction of the population is updated at each discrete time step. Such intermediate updated fractions are relevant in many biological situations. For instance, bacterial populations can comprise a fraction of slow-growing persister cells [20, 21, 22, 23]. Under stress, some bacteria can also undergo sporulation [24]. Such bet-hedging strategies, where a fraction of the population is dormant, facilitate population survival to harsh environmental conditions, and exist beyond bacteria. For instance, in annual plants, only a fraction of seeds may germinate each year, while others stay dormant [25, 26, 27, 28, 29, 30]. Besides, in a population where adults live for several years, with a reproduction season each year, and where mortality mainly arises from external factors such as infectious diseases or predation, the high mortality among young individuals results in an effective renewal of a fraction of the population during each reproductive season [31, 32]. Seed banks, where seeds are preserved and can yield offspring many generations later, can also give rise to complex partial generation overlaps [33, 34, 35].
To study this model, we focus on genetic drift and natural selection in a haploid population. We first assume that a fixed fraction of the population is updated at each discrete time step. Under Kimura’s diffusion approximation, we obtain a simple expression of the fixation probability of a mutant, and determine the average conditional and unconditional absorption times. In the neutral case, we further study the rate of approach to the steady state, and the coalescent process associated to our model. This allows us to determine the variance effective population size, the eigenvalue effective population size, and the coalescent effective population size, which coincide. Besides, the ancestral process converges to Kingman’s coalescent. Next, we generalize our model, first by taking into account fluctuations of the fraction that is updated at each discrete time step, and then by considering individuals with fluctuating lifetimes. This allows us to express the fixation probability as a function of the mean life time of individuals. Finally, we further generalize our model by incorporating selection on the lifetime as well as on the reproductive fitness, i.e. both on birth and death processes, assuming that lifetimes are geometrically distributed.
@@ II.1 Description of the model
Link to the Wright-Fisher model
@@
Link to the Moran model
@@ II.2 Mutant fixation probability
Let us consider a large-sized population $N\gg 1$ and assume that $s$ is of order $1/N$. These conditions allow us to employ Kimura’s diffusion approximation [36, 37, 38] (see Supplementary Material for mathematical details). The fixation probability $\phi(x)$ of the mutant type, starting from a fraction $x$ of mutants in the population, then satisfies the following Kolmogorov backward equation [36]:
@@
where $\Delta x$ is the variation of the mutant fraction occurring in the first discrete time step starting from the initial fraction $x$, and $\mathbb{E}(\Delta x)$ and $\mathbb{V}(\Delta x)$ denote respectively its expected value and variance. To leading order in $1/N$, they read (see Supplementary Material)
@@
This result generalizes over the fixation probability of both the Moran and the Wright-Fisher models in the diffusion regime. We recover them by taking $\rho=1/N$ (which is then neglected since we restrict to leading order in $1/N$), and $\rho=1$, respectively. In Fig. 2, we show the probability $\phi_{1}=\phi(1/N)$ of fixation in Eq. 11 for different values of the updated fraction $\rho$. We observe that intermediate values of $\rho$ lead to probabilities of fixation that are between those given by the Moran model and the Wright-Fisher model. The nonlinear dependence on $\rho$ of the fixation probability is illustrated in Figs. S1 and S2 of the Supplementary Material.
@@
In this regime, the mutant probability of fixation increases with the updated fraction $\rho$ at fixed $s$, forming a continuum between the Moran result ($\phi_{1}=s$) and the Wright-Fisher one ($\phi_{1}=2s$). We note that this is analogous to Haldane’s formula [39]
@@
where $\sigma^{2}=2-\rho$ is the variance of the number of offspring drawn from a Poisson distribution (see Supplementary Material).
Figure 3: The mean unconditional absorption time $t$, as well as the mean mutant fixation time $t^{*}$ and extinction time $t^{**}$ are shown in generations (i.e. rescaled with the generation time $1/\rho$) versus the initial mutant fraction $x$, for $N=10^{5}$. Each of these times is plotted for $\rho=1/N,0.25,0.5,0.75,1$, where increasing values of $\rho$ correspond to lighter colors. Different panels show different relative mutant fitness advantages $s$.
II.3 Mean absorption times
Kimura’s diffusion approximation also allows determining the mean absorption times, i.e. the time it takes for the mutant type to either fix in the population or to get extinct. Note that in the long term, fixation or extinction are the only possible outcomes: they are the two absorbing states of the system.
The mean unconditional absorption time $t(x)$ to reach either fixation or extinction, starting from an initial fraction $x$ of mutants with relative fitness $1+s$, satisfies [40]
$\frac{\mathbb{V}(\Delta x)}{2}\frac{\mbox{d}^{2}t}{\mbox{d}x^{2}}+\mathbb{E}(\Delta x)\frac{\mbox{d}t}{\mbox{d}x}=-1\,,$ (14)
and $t(0)=t(1)=0$. In the Supplementary Material, we provide integral solutions to this equation.
The mean absorption time $t^{*}$ conditioned on mutant fixation, and $t^{**}$ conditioned on mutant extinction, can be obtained in a similar way [40] (see Supplementary Material).
These three absorption times, rescaled to be expressed in generations, are shown in Fig. 3 versus the initial mutant fraction $x$ for different values of $\rho$ (different shades) and $s$ (different panels). We observe that increasing the updated fraction $\rho$ gradually increases all three absorption times, once they are expressed in generations. Besides, when the relative fitness advantage $s$ of the mutant decreases and goes toward effective neutrality, absorption times increase, and the unconditional one, $t$, becomes more symmetric around $x=1/2$, as for Moran and Wright-Fisher models. In addition, for $s>1/N$, the unconditional absorption time $t$ is close to the fixation time $t^{*}$ for sufficiently large $x$, as fixation then dominates. We further show in the Supplementary Material the symmetry property $t^{*}(1-x)=t^{**}(x)$.
II.4 Rate of approach to steady state in the neutral case
In this section, we consider neutral mutants, i.e. $s=0$. The Markov chain representing the time evolution of the number of mutants in the population possesses two absorbing states, namely fixation and extinction of the mutant type. Therefore, the transition matrix in our model, whose elements are given by Eq. 4, has two unit eigenvalues [40]. The next largest non-unit eigenvalue provides information on the rate of decrease of genetic diversity in the population. It is thus of interest to calculate it. Denoting by $X_{t}$ be the number of mutant at time $t$, we can show that (see Supplementary Material)
@@ II.4 Rate of approach to steady state in the neutral case
The convergence to steady state can be quantified by considering the vector $\bm{\pi}_{t}=(p_{0}(t),p_{1}(t),p_{2}(t),...,p_{N}(t))$ where $p_{k}(t)$ denotes the probability of having $k$ mutants at time $t$ starting from $1$ mutant at $t=0$. For $t\rightarrow\infty$, we have (see Supplementary Material)
@@
where ||.$||.||_{2}$ refers to the Euclidean norm and
$\bm{\pi}_{\infty}=\lim\limits_{t\rightarrow\infty}\bm{\pi}_{t}=(1-1/N,0,0,...,0,1/N).$ (19)
@@
In Fig. 4, this prediction is tested numerically for different updated fractions $\rho$ of the population. We find a good agreement between our simulation results and the long-time asymptotic behaviour given in Eq. 18. Increasing the updated fraction $\rho$ increases the time it takes to approach steady state, measured in generations. Again, our model leads to a continuum between the Moran model and the Wright-Fisher model.
@@
Figure 4: Approach to steady state in the neutral case. We show the Euclidean norm of $\bm{\pi}_{t}-\bm{\pi}_{\infty}$ (see main text) versus time in generations (i.e. rescaled with the generation time $1/\rho$), for $N=1000$ and with different updated fractions $\rho$. Full lines: simulation results obtained from $5.10^{7}$ runs. Dashed lines: asymptotic behaviour given in Eq. 18, with a multiplicative constant chosen for legibility.
The eigenvalue in Eq. 16 is related to the expected diversity $H(t)$ of the population (also called expected heterozygosity, especially in diploid populations), defined as the probability of drawing two individuals of different types (here, wild-type and mutant). Indeed, it reads:
$H(t)=\frac{\mathbb{E}\left[X_{t}(N-X_{t})\right]}{\binom{N}{2}}=\lambda_{*}^{t}H(0),$ (20)
where we have used the derivation of Eq. 15 (see Supplementary Material). Thus, $\lambda_{*}$ also characterizes the rate of loss of diversity in the population.
II.5 Coalescent
The coalescent traces the genealogy of a given sample of individuals backward in time to their most recent common ancestor. A large class of models converge to Kingman’s coalescent [41, 42, 43, 44] for neutral mutants in the limit of large population size. This includes Moran and Wright-Fisher models, but also the broad class of Cannings models [45, 46]. In the specific case of neutral mutants, our partial update model falls in the latter class. Therefore, under the proper time scaling, the coalescent process associated with our model in the neutral case also converges to Kingman’s coalescent in the limit of large populations [41, 46, 47], see Supplementary Material for details.
For completeness, let us also explicitly consider our model for finite population size $N$, still in the neutral case. A central quantity in coalescent processes is the probability $G_{i,j}$ that $i$ sampled individuals at discrete time step $t$ have exactly $j$ parents at $t-1$. In our model, we show in the Supplementary Material that this probability reads
[…]
$G_{i,i}$ $=1-\binom{i}{2}\frac{\rho(2-\rho)}{N}+o\left(\frac{\rho}{N}\right),$ (22)
$G_{i,i-1}$ $=\binom{i}{2}\frac{\rho(2-\rho)}{N}+o\left(\frac{\rho}{N}\right).$ (23)
Here too, taking $\rho=1/N$ and $\rho=1$ yields respectively the Moran model result and the Wright-Fisher model result [49]. These expressions also illustrate the link with Kingman’s coalescent (see Supplementary Material).
Furthermore, Eqs. 22 and 23 can be used to express the mean coalescence time, i.e. the mean time $T_{\text{MRCA}}$ to the most recent common ancestor of the population, for finite $N$ (see Supplementary Material). In the large-$N$ limit, the mean coalescence time reads [40, 50]
$T_{\text{MRCA}}=\frac{2(N-1)}{\rho(2-\rho)}.$ (24)
Note that this time is expressed in number of discrete time steps – thus, $\rho T_{\text{MRCA}}$ represents the coalescence time expressed in generations. The mean coalescence time $\rho T_{\text{MRCA}}$ is shown versus population size $N$ for various values of $\rho$ in Fig. S3 of the Supplementary Material, and the large-$N$ expression in Eq. 24 is observed to be a good approximation even for moderate $N$.
II.6 Effective population sizes
@@
The effective population size is the size a regular Wright-Fisher population should have to match a given property of the population considered. Several definitions can be found in the literature, each one matching a different property of the Wright-Fisher population, see [40, 51] for a review. Here, we determine several of these effective population sizes for a population of size $N$ within our partial update model.
Let us first consider the variance effective population size $N_{e,v}$. We can determine it in our model starting from the definition in Refs. [3, 40]:
$N_{e,v}\equiv\frac{x(1-x)}{\mathbb{V}_{g}(\Delta x)}=\frac{x(1-x)}{\mathbb{V}(\Delta x)/\rho}=\frac{N}{2-\rho}.$ (25)
Here, we denoted by $\mathbb{V}_{g}$ the variance accumulated over one generation [40]. Because the average generation time is $1/\rho$ in our model, $\mathbb{V}_{g}(\Delta x)$ is $1/\rho$ times the variance in Eq. 10, which allowed us to conclude.
@@
Next, let us consider the eigenvalue effective population size $N_{e,\lambda}$. Eq. 17 allows us to express it as [3]
$N_{e,\lambda}=\frac{1}{1-r}=\frac{N}{2-\rho}\,.$ (26)
Finally, since our model converges to Kingman’s coalescent for large $N$ in the neutral case with a proper time rescaling, we can define a coalescent effective size $N_{e,c}$ [52, 53, 54]. Note that this definition of the effective population size is sometimes considered more general than others, as a large variety of models converge to Kingman’s coalescent [41, 46]. Besides, the coalescent essentially contains all of the information that can be found in sampled genetic data [52, 53, 54]. We adopt the most general definition given in Ref. [54], which depends on the properties of the ancestral process, as well as on the time scaling (see Supplementary Material). The resulting effective coalescent size is
$N_{e,c}=\dfrac{N}{2-\rho}.$ (27)
The three definitions of effective population size coincide for our model, giving a unique expression $N_{e}$. Fig. S4 in the Supplementary Material illustrates how the effective population size $N_{e}$ increases non-linearly with the updated fraction $\rho$, from the Moran result, $N_{e}=N/2$, to the Wright-Fisher one, $N_{e}=N$. Thus, populations with partially-overlapping generations have intermediate effective population sizes.
@@ III.1 Fluctuating updated fraction
where $P_{i\rightarrow j}^{(N,M)}$ is given by Eq. 4. To obtain the mutant fixation probability in the diffusion approximation within this new model, we express the first two moments of $\Delta x$. The calculation is similar to the one presented in the Supplementary Material for the model with fixed updated fraction, and yields to leading order in $1/N$
@@
The first inequality follows from $\langle\rho^{2}\rangle\geq 0$. Moreover, since $\forall\rho\in\left[0,1\right]$, $\rho^{2}\leq\rho$, we have $\langle\rho^{2}\rangle\leq\langle\rho\rangle$ for any $\pi^{(N)}$, yielding the second inequality. This result entails that the fixation probabilities in this model are comprised between those of the Moran and of the Wright-Fisher models.
@@
$=\left[1-\langle\rho\rangle\right]^{k-1}\langle\rho\rangle\,.$ (34)
@@
$\langle T\rangle$ $=1/\langle\rho\rangle.$ (35)
@@ III.2 Fluctuating lifetimes
$\lim\limits_{t\rightarrow+\infty}p_{u}(t)=1/\langle T\rangle\,.$ (37)
@@
$\langle\rho\rangle=1/\langle T\rangle\,,$ (38)
@@
$\langle\rho^{2}\rangle/\langle\rho\rangle=1/\langle T\rangle\,.$ (39)
@@ III.3 Selection on the lifetime
Assuming as usual that $s$ is of order $1/N$, we obtain, to leading order in $1/N$ (see Supplementary Material):
@@ IV Discussion
We proposed a simple and tractable model that bridges the Wright-Fisher model and the Moran model, and can be seen as a particular case of the Chia and Watterson model [17]. We do not focus on the offspring distribution, but on the fraction $\rho$ of the population is updated at each discrete time step, using binomial sampling. This leads to a more tractable model where the Wright-Fisher and the Moran model emerge as simple limiting cases, respectively for $\rho=1$ and $\rho=1/N$. In our model, one generation corresponds to $1/\rho$ discrete time steps. We showed that the rescaled process, taking one unit of time as $N/\rho$ discrete steps, converges to a diffusion. Under the diffusion approximation, we obtained a simple expression of the mutant fixation probability, and expressed the conditional and unconditional mean absorption times. We also calculated the rate of approach to steady state. We further studied the genealogy of a population sample in the neutral case, which converges to Kingman’s coalescent for large population sizes [41, 45]. Convergence to the coalescent is achieved when taking one unit of coalescent time as $N/[\rho(2-\rho)]$ discrete time steps. This time scaling differs by a multiplicative constant from the one commonly used to show convergence to a diffusion process. We determined the variance, eigenvalue and coalescent effective population sizes in our model, and showed that they coincide. Throughout, we wrote key quantities characterizing the model as simple functions of the updated fraction $\rho$, and discussed its impact on these quantities.
@@
We extended our model to the case where the updated fraction $\rho$ fluctuates. We showed that the mean individual lifetime is the inverse of the mean updated fraction. We considered the case where the lifetime fluctuates. Finally, we examined the situation where selection occurs both on birth and on death, assuming that the lifetime of individuals is geometrically distributed. In all these cases, we obtained simple expressions of the fixation probability in the diffusion approximation.
Here, we did not model mutations, as we primarily focused on fixation probabilities and absorption times. Including mutations explicitly in our model would be an interesting extension. It would allow to derive the stationary distribution of mutants and wild-types in the population, since mutant fixation and extinction would no longer be absorbing states [40]. Previous works have investigated the impact of mutations on the effective coalescent size [54], as well as the diffusion approximation [57]. In the case of neutral mutants, our partial update model falls in the class of Cannings models [19, 18], which yield convergence to Kingman’s coalescent [41, 42, 45, 58] in the large-$N$ limit, even when parent-independent mutations are included. Furthermore, the stationary distribution of allele frequencies across generations is then the same as in the Wright-Fisher model [59].
Including fluctuations in the total population size in addition to fluctuations of the updated fraction would account for “sweepstakes effect” observed for marine organisms, where adults release a large number of gametes. While most gametes are lost to external perturbations, reproduction is a chance event that can produce a very large number of offspring [60, 61].
@@
Here, we assumed discrete time steps, as in most classic models of population genetics. This is relevant in many situations, e.g. annual reproduction. Moreover, when continuous time is more relevant, as in bacterial reproduction, mutant replications remain synchronized for some time, as they stem from the same ancestor [62]. Besides, while we focused on haploids, our model could also be extended to diploids, and various hypotheses on reproduction could be investigated.
Finally, we focused on a well-mixed population with fixed size. It would be interesting to extend our model to spatially structured populations [8, 63, 11, 64], or to fluctuations of the population size [65].
Acknowledgments
A. A. thanks Félix Foutel-Rodier for useful suggestions. This research was partly funded by the Swiss National Science Foundation (SNSF) (grant No. 315230_208196, to A.-F. B.), by the Chan-Zuckerberg Initiative (CZI, to A.-F. B.), and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851173, to A.-F. B.). C. L. acknowledges funding by the Agence Nationale de la Recherche (grants ANR-20-CE30-0001 and ANR-21-CE45-0015, to C. L.).
Supplementary Material